The formula to calculate the real interest rate (i) is:

i= (if + f)/ (1+f)

Question 3 options:

True
False
Answer #1

Answer: False

Reason: the formula for inflated interest rate if = i + f + if [ where i= real interest rate, f= inflation rate and

if = (real interest rate * inflation rate)]

thus

i + f + if = if

or, i+if = if - f

or, i(1+f) = if - f

or, i = (if - f)/ (1+f)

This formula dose not match with the formula given in the question thus the answer is false.